How does x+6equal 7-x
Answer by stanbon(75887)

How does x+6equal 7-x
---
Solve x+6 = 7-x
Add x to both sides.
Subtract 6 from both sides:
2x = 1
Divide by 2 to get:
x = 1/2
=============
Cheers,
Stan H.
=============